How do I graph the following line? Through (-4,-1) with m=3. So far I have gotten- Y-y1=m(X-x1) Y--1=3(X--4) Y+1=3x+12 -3x+y=12-1 -3x+y=11 y=3x+11. This is where I get stuck.
Answer by stanbon(75887)

How do I graph the following line? Through (-4,-1) with m=3.
y-3 = m(x+4)
y = 3(x+4)+3
y = 3x+12 + 3
y = 3x + 15
==========
Cheers,
Stan H.